OpenStudy (anonymous):

−2(x + 3) = −2(x + 1) − 4

sam (.sam.):

uh cannot be defined

sam (.sam.):

-2(x+3)=-2(x+1)-4 Multiply -2 by each term inside the parentheses. -2(x+3)=-2x-2-4 Subtract 4 from -2 to get -6. -2(x+3)=-2x-6 Multiply -2 by each term inside the parentheses. -2x-6=-2x-6 Move all terms not containing x to the right-hand side of the equation. -6=-6 Since -6=-6, the equation will always be true. Always True

OpenStudy (anonymous):

Would it be all real numbers? or no solution?